OpenStudy (anonymous):

complete each ordered pair so that it is a solution of the given linear equation. x-3y=-4; ( ,5) , (-1, ) the first ordered pair is ( ,5 )

OpenStudy (academicgurusinc):

Answer: (11, 5), (-1, 1) Solution: If ( ,5) = (x, 5) and (-1, ) = (-1, y) then: for (,5): Rewrite x - 3y=-4 as x -3 (5) = -4 Solve for x: x-3(5)= -4 x-15 = -4 x=-4 + 15 x= 11 for (-1,y) Rewrite x - 3y = -4 as -1- 3y = -4 Solve for y: -1-3y=-4 -3y= -4+ 1 -3y =-3 y= 1 Any further queries feel free to contact us at info@academicgurusinc.com We hope we were able to help. Follow us on facebook and twitter.